40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 30 13:14:03 2008, Koji, Summary, IOO, History of the MC abs length MC_length.png
    Reply  Wed Jul 30 15:12:08 2008, rana, Summary, IOO, History of the MC abs length 
Message ID: 768     Entry time: Wed Jul 30 13:14:03 2008     Reply to this: 770
Author: Koji 
Type: Summary 
Category: IOO 
Subject: History of the MC abs length 
I was notified by Rob and Rana that there were many measurements of the MC abs length (i.e. modulation 
frequencies for the IFO.) between 2002 and now.

So, I dig the new and old e-logs and collected the measured values of the MC length, as shown below. 

I checked the presence of the vent for two big steps in the MC length. Each actually has a vent. 
The elog said that the tilt of the table was changed at the OMC installation in 2006 Oct.
It is told that the MC mirrors were moved a lot during the vent in 2007 Nov.

Note:
o The current modulation freq setting is the highest ever.
o Rob commented that the Marconi may drift in a long time.
o Apparently we need another measurement as we had the big earthquake.

My curiosity is now satified so far.

Local Time	3xFSR[MHz]	5xFSR[MHz]	MC round trip[m]	Measured by
----------------------------------------------------------------------------
2002/09/12	33.195400 	165.977000 	27.09343		Osamu
2002/10/16	33.194871 	165.974355 	27.09387		Osamu
2003/10/10	33.194929 	165.974645 	27.09382		Osamu
2004/12/14	33.194609 	165.973045 	27.09408		Osamu
2005/02/11	33.195123 	165.975615 	27.09366		Osamu
2005/02/14	33.195152 	165.975760 	27.09364		Osamu
2006/08/08	33.194700 	165.973500 	27.09401		Sam
2006/09/07	33.194490 	165.972450 	27.09418		Sam/Rana
2006/09/08	33.194550 	165.972750 	27.09413		Sam/Rana
----2006/10 VENT OMC installation	
2006/10/26	33.192985 	165.964925 	27.09541		Kirk/Sam
2006/10/27	33.192955 	165.964775 	27.09543		Kirk/Sam
2007/01/17	33.192833 	165.964165 	27.09553		Tobin/Kirk
2007/08/29	33.192120 	165.960600 	27.09611		Keita/Andrey/Rana
----2007/11 VENT Cleaning of the MC mirrors
2007/11/06	33.195439 	165.977195 	27.09340		Rob/Tobin
2008/07/29	33.196629 	165.983145 	27.09243		Rob/Yoichi
Attachment 1: MC_length.png  29 kB  Uploaded Wed Jul 30 14:29:37 2008  | Hide | Hide all
MC_length.png
ELOG V3.1.3-